A particle of charge q = 40 µC is placed at a corner of a Gaussian cube of edge a = 9.8 cm. What is the flux through (a) each cube face forming that corner and (b) each of the other cube faces?
Answer #1

a) 0.875*40*(10^-6)/(3*8.85*10^-12) = 1.318*10^6 Vm

b) 0.125*40*(10^-6)/(3*8.85*10^-12) = 1.883*10^6 Vm